When 25.00mL of sodium hydroxide solution of unknown concentration was titrated against a standard hydrochloric acid solution of concentration 0.098 mol L-1, 14.73 mL of the acid was needed to completely neutralize the base. Calculate the concentration of the sodium hydroxide solution
1
Expert's answer
2019-05-02T07:30:35-0400
"NaOH +HCl\rightarrow NaCl + H_2O"

Find moles of HCl:


"c= \frac{n}{V}"

"\therefore n= c\times V"

"n(HCl)= 0.098\frac{mol}{L}\times14.73\times10^{-3}L=1.444\times10^{-3} mol"

According to equation "n(HCl):n(NaOH)=1:1" , then "n(NaOH)=n(HCl)=1.444\times10^{-3} mol"

Find concentration of NaOH solution:


"c= \frac{n}{V}= \frac{1.444\times10^{-3}mol}{25.00\times 10^{-3}L}=0.058\frac{mol}{L}"